# Detecting Spike Wave Discharges (SWD) using 1-dimensional Residual UNet

### v1

@@ Abstract
The manual labeling of events in electroencephalography (EEG) records is time-consuming. This is especially true when EEG recordings are taken continuously over weeks to months. Therefore, a method to automatically label pertinent EEG events reduces the manual workload. Spike wave discharges (SWD), which are the electrographic hallmark of absence seizures, are EEG events that are often labeled manually. While some previous studies have utilized machine learning to automatically segment and classify EEG signals like SWDs, they can be improved. Here we compare the performance of 14 machine learning classifiers on our own manually annotated dataset of 961 hours of EEG recordings from C3H/HeJ mice, including 22,637 labeled SWDs. We find that a 1D UNet performs best for labeling SWDs in this dataset. We also improve the 1D UNet by augmenting our training data and determine that scaling showed the greatest benefit of all augmentation procedures applied. We then compare the 1D UNet with data augmentation, AugUNet1D, against a recently published time- and frequency-based algorithmic approach called "Twin Peaks". AugUNet1D showed superior performance and detected events with more similar features to the SWDs labeled manually. AugUNet1D, pretrained on our manually annotated data or untrained, is made public for others users.
@@ Introduction
Absence seizures are brief episodes of altered consciousness characterized by sudden staring spells, typically lasting 5-20 seconds[1, 2]. Individuals often lose consciousness before returning to normal awareness with no memory of the event. This epilepsy is more prevalent in children than adults and can significantly impact academic performance and social development, as they can occur dozens of times a day, causing frequent lapses in attention[3, 4]. Understanding the underlying neural mechanisms will help develop more targeted treatments and improve diagnostic precision.
@@
EEG (electroencephalography) is the gold standard for detecting absence seizures due to its ability to capture the characteristic electrical brain activity patterns that occur during these episodes. During absence seizures in humans, EEG recordings show a distinctive pattern of generalized 3-Hz spike-wave discharges (SWD) that appear suddenly across much of the brain nearly simultaneously. This pattern is so specific to absence seizures that it serves as a diagnostic hallmark, distinguishing absence epilepsy from other types of epilepsy and attention disorders[1, 2]. The EEG is particularly valuable because absence seizures often occur without obvious external symptoms.
@@
Long time series data, like those collected during prolonged continuous EEG recordings, can be tedious and time consuming to label manually. Machine learning and neural network approaches offer opportunities for automating this process. In fact, while these approaches been applied to similar EEG-labelling use cases for decades[5], innovations over the last several years have culminated in end-to-end solutions that automatically label events in raw or minimally preprocessed EEG signals [6, 7, 8]. These approaches extract hierarchical representations from the raw data rather than rely on features extracted from time, frequency, and time-frequency domains [9, 10, 11].
@@
Hybrid convolutional neural network-long short-term memory (CNN-LSTM) architectures have shown particular promise by combining spatial feature extraction from CNNs with temporal modeling capabilities of LSTMs [12, 13]. More recently, specialized architectures designed specifically for EEG analysis (e.g., EEGNet [14]), time series classification (e.g., InceptionTime [15]), and dense segmentation tasks (e.g., UNet-based models [16, 17]) have emerged as promising approaches for automated physiological signal analysis. Additionally, transformer-based architectures adapted from computer vision, such as DETRtime [18] for temporal event detection and SalientSleepNet [19] for sleep stage classification, represent the latest trend in applying attention mechanisms to EEG time series analysis.
@@
With this in mind, we compared the performance 14 machine learning classifiers to determine which is superior for the detection of SWDs in a mouse model of absence epilepsy, the C3H/HeJ mouse strain [20, 21, 22]. After determining that the 1D residual UNet performed best, we further improved performance by applying augmentation procedures to the training data. The final product, AugUNet1D, consistently performed well across recordings from 10 mice in our test dataset. We also compared AugUNet1D to a recent algorithmic method using time- and frequency-based features, "Twin Peaks" [23]. AugUNet1D again showed superior performance, detecting events with features similar to manually labeled SWDs.
Overall, results indicate that AugUNet1D is an effective and reliable method for labeling SWD in EEG from C3H/HeJ mice. We provide code for using a version of AugUNet1D that has been pre-trained on our manually annotated dataset, which will be useful for those recording SWDs in C3H/HeJ mice. We also provide code for a naive version of the network that users can train on their own manually labeled data. The untrained version of AugUNet1D may be applied to and effective for data from other organisms, different types of seizures, or other EEG events altogether.
@@
Figure 1: Schematic of EEG data preprocessing, augmentation, and the architecture of the AugUNet1D network.
@@
Animals were anesthetized with vaporized isoflurane (2-4% for induction, 1-2% for maintenance) mixed with pure oxygen and delivered at a rate of 1 liter/min. The scalp was first shaved and then depilatory cream was applied to remove all remaining hair. The scalp was then cleaned with alternating scrubs of betadine and 70% ethanol three times. A sagittal incision was made starting between the eyes and ending between the ears. The underlying fascia was scrubbed away vigorously with sterile cotton tipped applicators and an applicator dipped in hydrogen peroxide (3%) was used to clean the skull and increase the visibility of cranial sutures. Burr holes were made above the secondary motor or primary somatosensory cortex unilaterally or bilaterally in some cases depending on the experiment design. An additional burr hole was made in the center of the interparietal bone overlying the cerebellum. A custom electrode array, consisting of insulated stainless steel wires (A-M systems part no. 791400) with approximately 500µm insulation stripped from the end soldered to a 6-position header of connector pins (Digikey part no. 1212-1786-ND), was then lowered into the surgical field. The wires were carefully placed under the skull in the burr holes using forceps, one at a time, and UV-curing glue (SDI Wave) was applied and cured to keep the wires in place. When all wires were positioned and glued, the header of connector pins was cemented to the remaining exposed skull with dental cement (Metabond, Parkell part no. S380). Skin at the perimeter of the incision was pulled tightly around the cement base and adhered to it with cyanoacrylate glue or sutured together at the posterior end of the incision. Ketoprofen (5mg/kg) was administered peri-operatively to reduce pain and inflammation. Animals were given at least 7 days to recover before any other procedures were conducted, including EEG recording.
@@
In all experiments, EEG signals were first passed through an operational amplifier to buffer the small voltage fluctuations before subsequent amplification. These op-amps were integrated into custom-made PCBs which connected to the header of pins implanted on the head. In some cohorts, signals were then further amplified 10,000 times by a differential amplifier (A-M systems model 3500) and digitally sampled at 1kHz (AD Instruments Powerlab 16/35 or PowerLab C using LabChart). In another a separate cohort, data were acquired with a different EEG monitoring system (Grass AURA-64 LTM using TWin software) sampling at 400Hz. In all experiments, the cerebellar electrode served as the common reference for other recording channels. Recordings were continuous, 24 hours per day, and were stopped, saved, and restarted once per day. All data analyzed in these experiments was recorded from the electrode in the secondary motor cortex.
@@
Spike-wave discharges were labeled manually by three trained observers. All observers used the following minimum criteria for labeling SWDs: the event must comprise at least 5 clear rhythmic spike-wave complexes occurring at short, regular intervals separated by at least 50 milliseconds. The beginning of the events was defined as the time immediately preceding the initial negative inflection of the first spike in the spike-wave discharge. The end was defined as the time after the minimum value of the last spike occurred. Manual labeling was performed in Clampfit software (Molecular Devices San Jose, CA) by placing the first and second cursors such that the entire event fell between them. All observers timed their own manual labeling sessions and the time was subsequently saved on a spreadsheet.
Figure 2: Example of a seizure in an EEG trace. Note the high amplitude, rhythmic events that define SWD occurring during the correctly segmented orange portion of the trace.
The training data consisted of 5 days of continuous EEG recording in 8 mice, totaling 961.3 hours of recording. Manual scoring took a total of 108.09 hours across all three experimenters, resulting in a ratio of 8.89 data hours scored per hour of scorer time (Table 1).
Scorer Manual Scoring Hours Data Hours Time scored to time spent scoring ratio
@@
Table 1: Time spent scoring the manually labeled dataset
In total, across all 5 days and 8 mice, 22,637 SWDs were manually identified (Figure 3B). The average SWD rate across these recordings was 23.55 ± 9.3 per hour (mean ± SD, N = 8 mice). We also calculated the percentage of total SWDs that were recorded from each mouse and found it ranged from 1115 SWDs (4.9%) to 4570 SWDs (20.2%). The average duration across all events was 5.83 ± 2.37 seconds and varied little between mice (Figure 3A). We also characterized the spectral composition of all manually labeled SWD events (Figure 3C). To this end, the power spectral density of each manually labeled SWD was calculated and the peak frequency was determined (Figure 3C, bottom). The mean peak frequency of all SWDs was 5.72 ± 0.75Hz, again showing little variation across mice (Figure 3D). The observed event rates, duration, and spectral features in our training dataset are very similar to those in previous reports [20, 21, 22].
@@
Classifying Sleep and Noise in EEG Records
Noise and sleep epochs were identified in all recordings in the test dataset. To identify periods contaminated with noise artifacts caused by cable swinging, poor connection, and/or head banging, EEG traces were first segmented into consecutive 20-second blocks. The average value of the EEG was then computed for each of these blocks. Any blocks that had an average 20 standard deviations over the mean EEG were classified as noise epochs. Sleep epochs were defined as periods of 20 seconds or more when the delta-filtered (0.1 to 4Hz) power envelope maintained an amplitude above a minimum detection threshold. Additionally, only periods that also crossed a second, higher threshold at least once during that time were considered sleep epochs. These two thresholds were determined by first estimating a distribution of the delta power envelope across an entire recording, then finding the 2 largest peaks in this distribution. The lower-value peak corresponds to periods delta power during wakefulness. The higher-value peak corresponds to delta during deep slow wave sleep and was therefore used as the higher threshold. The midpoint value between the two peaks was used as the lower, minimum detection threshold when defining sleep epochs.
@@
The training data collected was recorded at a 1000Hz while all our labeled testing data was recorded at 400Hz. To take into account variable sampling rates, we used torchaudio package from PyTorch project to resample all input data to a standard sampling rate of 100Hz. ‘sinc_interp_hann’ is a PyTorch function that performs sinc interpolation with Hann windowing for resampling audio or other time-series signals. This function implements a high-quality interpolation method that uses the sinc function $(sin(\pi x)/(\pi x))$ as the basis for reconstruction, which is theoretically optimal for bandlimited signals according to the Nyquist-Shannon sampling theorem. The addition of Hann windowing helps reduce spectral leakage and ringing artifacts that can occur with pure sinc interpolation.
In the context of EEG signal processing for our use case, ‘sinc_interp_hann’ is useful for standardizing sampling rates across different recording devices or datasets, ensuring that neural network models receive consistently sampled input data. The high-quality interpolation is especially important for EEG analysis because it preserves the precise timing and frequency characteristics of brain signals that are critical for accurate seizure detection or other neurological event classification.
Therefore, we take our EEG signal and ground truth labels $x[n]$ and resample using $\mathcal{R}_{f_{s}\rightarrow f_{t}}:\mathbb{R}^{N_{s}}\rightarrow\mathbb{R}^{N_{t}}$
Where $\mathcal{R}_{f_{s}\rightarrow f_{t}}$ is the torchaudio Resample transform that converts an EEG signal from source sampling rate $f_{s}$ to target sampling rate $f_{t}$.
@@
For our seizure detection task, we adopted a residual 1D U-Net architecture. The U-Net was originally developed by Ronneberger et al. (2015) [16] for biomedical image segmentation, where it demonstrated remarkable success in precisely delineating cellular structures and tissue boundaries in microscopy images. The architecture’s key innovation is its symmetric encoder-decoder structure with skip connections: the encoder progressively downsamples the input to capture hierarchical spatial features and contextual information, while the decoder upsamples to recover fine-grained spatial details. Crucially, skip connections bridge corresponding encoder and decoder layers, allowing the network to combine high-level semantic features with low-level spatial precision which is essential for accurate pixel-wise segmentation.
@@
The integration of residual connections with U-Net architectures has a well-established history in computer vision, where residual U-Net variants have demonstrated substantial improvements over vanilla U-Net. Zhang et al. (2018) [26] introduced ResUNet for road extraction from aerial images, while Alom et al. (2018) [27] proposed the Recurrent Residual U-Net (R2U-Net) for medical image segmentation. Building on these successes, we incorporated residual connections within each encoding and decoding block of our 1D U-Net architecture. These shortcut connections, inspired by ResNet [28] architectures, allow gradients to flow directly through the network during backpropagation, addressing the vanishing gradient problem and enabling stable training of deeper networks.

### v2

# Combining Residual U-Net and Data Augmentation for Dense Temporal Segmentation of Spike Wave Discharges in Single-Channel EEG

@@ Abstract
Manual annotation of spike-wave discharges (SWDs), the electrographic hallmark of absence seizures, is labor-intensive for long-term electroencephalography (EEG) monitoring studies. While machine learning approaches show promise for automated detection, they often struggle with cross-subject generalization due to high inter-individual variability in seizure morphology and signal characteristics. In this study we compare the performance of 15 machine learning classifiers on our own manually annotated dataset of 961 hours of EEG recordings from C3H/HeJ mice, including 22,637 labeled SWDs and find that a 1D U-Net performs the best. We then improve its performance by employing residual connections and data augmentation strategies combining amplitude scaling, Gaussian noise injection, and signal inversion during training to enhance cross-subject generalization. We also compare our method, named AugUNet1D, to a recently published time- and frequency-based algorithmic approach called "Twin Peaks" and show that AugUNet1D performs better on our dataset. AugUNet1D, pretrained on our manually annotated data or untrained, is made public for other users.
@@ Introduction
Childhood Absence Epilepsy (CAE) is characterized by prominent electrographic signals and concomitant loss of consciousness [1, 2]. This pediatric form of epilepsy can significantly impact academic performance and social development, as seizures often occur dozens of times a day, causing frequent lapses in attention [3, 4]. Understanding the underlying neural mechanisms of CAE will advance the development of more targeted treatments and improve diagnostic precision. Clearly, accurate and efficient seizure detection is a necessary component for such development.
@@
Electroencephalography (EEG) remains essential for detecting absence seizures. EEG recordings of absence seizures reveal a distinctive pattern of generalized 3-Hz spike-wave discharges (SWDs) that appear suddenly across much of the brain nearly simultaneously. The SWD serves as a diagnostic hallmark that distinguishes absence epilepsy from other types of epilepsy and attention disorders [1, 2]. Indeed, as absence seizures often occur without obvious external symptoms, EEG recordings are necessary for proper diagnosis.
@@
Manual evaluation of EEG recordings remains the gold standard for seizure phenotype assessment, largely due to the inadequacy of current automatic detection algorithms. Machine learning and neural network approaches offer opportunities for automating this process. In fact, while these approaches have existed for decades[5], innovations over the last several years have culminated in solutions that automatically label events in raw or minimally preprocessed EEG signals [6, 7, 8]. These approaches extract hierarchical representations from the raw data rather than rely on features extracted from time, frequency, and time-frequency domains [9, 10, 11]. However, such approaches have not yet been tested for the specific use case of automatically detecting SWDs.
@@
Hybrid convolutional neural network-long short-term memory (CNN-LSTM) architectures have shown particular promise by combining spatial feature extraction from CNNs with temporal modeling capabilities of LSTMs [12, 13]. More recently, specialized architectures designed specifically for EEG analysis (e.g., EEGNet[14]), time series classification (e.g., InceptionTime[15]), and dense segmentation tasks (e.g., UNet-based models[16, 17]) have emerged as promising approaches for automated physiological signal analysis. Additionally, transformer-based architectures adapted from computer vision, such as DETRtime[18] for temporal event detection and SalientSleepNet[19] for sleep stage classification, have also been applied to EEG time series analysis.
@@
Considering these numerous approaches to automated event detection, here, we compared the performance of 15 machine learning classifiers to identify optimal SWD detection strategies in the C3H/HeJ mouse model of absence epilepsy [20, 21, 22]. After determining that the 1D residual U-Net performed best, we further improved performance by applying augmentation procedures to the training data. The final result, AugUNet1D, consistently performed well across recordings from 10 mice in our test dataset. We also compared AugUNet1D to a recent algorithmic method using time- and frequency-based features called "Twin Peaks"[23]. AugUNet1D again showed superior performance, detecting events with features similar to manually labeled SWDs. Overall, our results demonstrate that AugUNet1D is an effective and reliable method for labeling SWDs in EEG from C3H/HeJ mice. Moreover, we establish a workflow for utilizing an untrained version of AugUNet1D that can be applied to detect a diverse set of electrographic events across different biological systems.
@@
Figure 1: Schematic of EEG data preprocessing, augmentation, and the architecture of the AugUNet1D network. We resample all signals to 100Hz, use augmentations like scaling, inversion and Gaussian noise and apply them randomly to the training signal, and then feed them into a residual 1D-UNet that predicts labels at every time point. We minimize Dice Loss using an Adam optimizer.
@@
Animals were anesthetized with vaporized isoflurane (2-4% for induction, 1-2% for maintenance) mixed with pure oxygen and delivered at a rate of 1 liter/min. The scalp was first shaved and then depilatory cream was applied to remove all remaining hair. The scalp was then cleaned with alternating scrubs of betadine and 70% ethanol three times. After creating a sagittal incision, the underlying fascia was scrubbed away with sterile cotton tipped applicators and an applicator dipped in hydrogen peroxide (3%) was used to clean the skull and increase the visibility of cranial sutures. Burr holes were made above the secondary motor or primary somatosensory cortex. An additional burr hole was made in the center of the interparietal bone overlying the cerebellum. A custom electrode array, consisting of insulated stainless steel wires (A-M systems part no. 791400) with approximately 500µm insulation stripped from the end soldered to a 6-position header of connector pins (Digikey part no. 1212-1786-ND), was then lowered into the surgical field. The wires were carefully placed under the skull in the burr holes using forceps and set into place with UV-curing glue (SDI Wave). When all wires were positioned and glued, the electrode array was cemented to the remaining exposed skull with dental cement (Metabond, Parkell part no. S380). Skin at the perimeter of the incision was pulled tightly around the cement base and secured with cyanoacrylate adhesive or sutured together at the posterior end of the incision. Ketoprofen (5mg/kg) was administered peri-operatively to reduce pain and inflammation. Animals were given at least seven days to recover prior to EEG recording.
@@
EEG signals were first passed through an operational amplifier to buffer small voltage fluctuations before subsequent amplification. These amplifiers were integrated into custom-made PCBs which connected to the header of pins implanted on the head. In some cohorts, signals were then further amplified 10,000 times by a differential amplifier (A-M systems model 3500) and digitally sampled at 1kHz (AD Instruments Powerlab 16/35 or PowerLab C using LabChart) or 400Hz (Grass AURA-64 LTM using TWin software). In all experiments, the cerebellar electrode served as the common reference for other recording channels. Recordings were continuous–24 hours per day–and were stopped, saved, and restarted once per day. All data analyzed in these experiments was recorded from the electrode in the secondary motor cortex. After the final recordings session, animals were deeply anesthetized with a lethal intraperitoneal injection of 60mg/kg Euthasol (pentobarbital sodium and phenytoin sodium) and then, after the loss of pedal reflex was confirmed, transcardially perfused with phosphate-buffered saline followed by 4% paraformaldehyde.
@@
SWDs were labeled manually by three trained researchers. The following criteria for labeling SWDs were used: (1) the electrographic event must include at least 5 clear rhythmic spike-wave complexes; and (2) the complexes must occur at short, regular intervals separated by at least 50 milliseconds. Manual labeling was performed in Clampfit (Molecular Devices, San Jose, CA) by placing cursors to bracket the entire SWD (Figure 2). The duration of manual labeling sessions was recorded.
Figure 2: Examples of SWDs in raw EEG traces. A) An example EEG trace showing 3 SWDs, each indicated by gray horizontal lines that correspond to manual labels. B) Expanded view of the portion of the trace within the dotted box in A. High amplitude, rhythmic spike-and-wave patterns that define SWD are clearly visible.
The training data consisted of five days of continuous EEG recording in eight mice, totaling 961.3 hours of recording. On average, it required one hour of manual work to label 8.89 hours of EEG data (Table 1).
Scorer Duration to Label (Hours) Duration of Data (Hours) Data-To-Labeling Ratio
@@
Table 1: Duration (hours) spent scoring the manually labeled dataset
In total, 22,637 SWDs were manually labeled across five recording days in eight mice (Figure 3A). SWD occurrence was variable across mice, ranging from 1,115 (4.9%) to 4,570 (20.2%) SWDs per mouse. The average SWD rate across all mice was 23.55 ± 9.3 per hour (mean ± SD, N = 8 mice). The average duration across all SWDs was 5.83 ± 2.37 seconds and varied little between mice (Figure 3B). We also characterized the spectral composition of all manually labeled SWDs (Figure 3C). To this end, the power spectral density of each manually labeled SWD was calculated, from which the peak frequency was determined (Figure 3C, bottom). The mean peak frequency of all SWDs was 5.72 ± 0.75Hz, again showing little variation across mice (Figure 3D). The observed event rates, duration, and spectral features in our training dataset are very similar to those in previous reports [20, 21, 22]. Thus, SWD features are generally preserved in C3H/HeJ mice, despite variable SWD occurrence across individuals.
@@
Classifying Noise and Sleep in EEG Records
Because EEG power broadly increases during periods of mechanical noise and during sleep, automated methods often falsely detect events during these periods. To assess how various methods perform during such conditions, noise and sleep epochs must be identified in our recordings. To define noise epochs, first, the mean and standard deviation of EEG amplitudes across the entire recording are computed. Then, traces are segmented into consecutive 5-second blocks. Blocks containing amplitudes 20 standard deviations or more beyond the mean are classified as noise (Figure 4A and B, yellow shading).
Sleep epochs are defined as periods with elevated power in the 0.1-4Hz frequency band (Figure 4C and 4D, blue shading). Thus, to identify sleep, the instantaneous amplitude (Hilbert envelope) of the 0.1-4Hz-band EEG is first computed as the magnitude of the analytic signal. Then, a distribution of the instantaneous amplitudes is estimated (Figure 4E). This distribution is bimodal: the lower-amplitude peak corresponds to wakefulness. The higher-amplitude peak corresponds to slow wave sleep. The midpoint between the two peaks is used as the lower threshold (red dotted line in Figures 4D and 4E) to detect the on- and off-set of sleep. Any epoch that does not cross a secondary threshold (the higher-value peak; blue dotted line in Figures 4D and 4E) is not considered sleep. Sleep epochs occurring within 20 seconds of each other are merged.
Figure 4: Examples of noise and sleep epochs in EEG traces. A) EEG trace showing 6 discrete noise epochs identified by yellow shading. B) Trace shows the maximum absolute EEG amplitude within consecutive 5-second windows. The orange dotted horizontal line is the threshold used to detect noise and represents 20 standard deviations above the overall absolute mean EEG amplitude. C) Representative EEG trace with sleep epochs indicated by blue shading. D) The instantaneous amplitude derived from the Hilbert envelope of the 0.1-4Hz-filtered EEG. The red dotted horizontal line indicates the minimum detection threshold that determines onset and offset of sleep epochs. The trace must also cross the blue dotted horizontal line at least once during an epoch to be classified as sleep. E) Distribution of the amplitude from the Hilbert envelope of the 0.1-4Hz-filtered EEG. This distribution is generated from the entire recording and therefore contains data recorded when the mouse is awake (low amplitude peak) and asleep (high amplitude peak). Horizontal gray bars in A and C indicate manually labeled SWDs.
@@
To account for variable sampling rates, we used the torchaudio package established by the PyTorch project to resample all input data to a standard sampling rate of 100Hz. ‘sinc_interp_hann’ is a PyTorch function that performs sinc interpolation with Hann windowing for resampling audio or other time-series signals.
In the context of EEG signal processing for our use case, ‘sinc_interp_hann’ is useful for standardizing sampling rates across different recording devices or datasets, ensuring that neural network models receive consistently sampled input data. The associated high-quality interpolation is especially important for EEG analysis because it preserves the precise timing and frequency characteristics of brain signals that are critical for accurate seizure detection or other neurological event classification strategies.
Therefore, our EEG signal and ground truth labels $x[n]$ are resampled using $\mathcal{R}_{f_{s}\rightarrow f_{t}}:\mathbb{R}^{N_{s}}\rightarrow\mathbb{R}^{N_{t}}$
Where $\mathcal{R}_{f_{s}\rightarrow f_{t}}$ is the torchaudio Resample transform that converts an EEG signal from its source sampling rate $f_{s}$ to its target sampling rate: $f_{t}$
@@
For our seizure detection task, we adopted a residual 1D U-Net architecture. The U-Net was originally developed by Ronneberger et al. (2015) [16] for biomedical image segmentation, where it demonstrated remarkable success in precisely delineating cellular structures and tissue boundaries in microscopy images. The architecture’s key innovation is its symmetric encoder-decoder structure with skip connections: the encoder progressively downsamples the input to capture hierarchical spatial features and contextual information, while the decoder upsamples to recover fine-grained spatial details. Crucially, skip connections bridge corresponding encoder and decoder layers, allowing the network to combine high-level semantic features with low-level spatial precision, an essential property for accurate pixel-wise segmentation.
@@
The integration of residual connections with U-Net architectures has a well-established history in computer vision, where residual U-Net variants have demonstrated substantial improvements over vanilla U-Net. Zhang et al. (2018) [26] introduced ResUNet for road extraction from aerial images, whereas Alom et al. (2018) [27] proposed the Recurrent Residual U-Net (R2U-Net) for medical image segmentation. Building on these successes, we incorporated residual connections within each encoding and decoding block of our 1D U-Net architecture. These shortcut connections, inspired by ResNet [28] architectures, allow gradients to flow directly through the network during backpropagation, addressing the vanishing gradient problem and enabling stable training of deeper networks.